A 500g lump of clay is flying horizontally at 2.5 m/s when it hits a stationary vertical uniform rod, hitting it at 25 cm below the top. The bar is 0.75m long, has a mass of 2.2 kg, and is hinged at its base. The lump of clay sticks to the bar. What is the magnitude of the angular velocity of the bar just after the collision?

A. 5 rad/s

B. 1.2 rad/s

C.0.7 rad/s

D.1.5 rad/s

E. 2.7 rad/s

The answer is B. 1.2 rad/s, however I need to know how to get to this answer!

Thanks! :)
Answer #1

using conservation of angular momentum

mvr + (1/3)*M L^2 * 0 = ( 1/3 ML^2 + mr^2) w

0.5* 2.5* 0.5 = ( 1/ 3* 2.2* 0.75^2 + 0.5* 0.5^2) w

w = 1.16≈ 1.2 rad/s

======

Comment in case any doubt, will reply for sure.. Goodluck